What is the mass in grams of 1.90x10^24 atoms of scandium
1
Expert's answer
2018-03-10T03:54:07-0500
m (Sc) = n×M, where n is amount of substance, mole; M is molar mass, g/ mole
n (Sc) = N/Na, N is number of atoms, Na is Avogadro constant
n = 1.90×10^24/6.02×10^23 = 3.2 mole
m = 3.2 × 45 = 144 g